OpenStudy (anonymous):

please help me solve this problem. Simplify 64w^3x^2 divided by 72wx

OpenStudy (anonymous):

\[\frac{64w^{3}x^{2}}{72wx}\] it might be easier to see it like this. If you notice the variables on the numerator and denominator, you'll see that you have similar variables. What that means is that you can reduce the variables directly. \[\frac{w^{3}}{w} = w^{2}\] \[\frac{x^{2}}{x} = x\] you can also break down the 64 and 72 into it's prime numbers to help cancel out terms. \[64 = 2*2*2*2*2*2\] \[72 = 2*2*2*3*3\] so if we represent it like this \[\frac{2*2*2*2*2*2}{2*2*2*3*3}\] we can cross out three of the 2s and be left with \[\frac{2*2*2}{3*3}=\frac{8}{9}\] comebine everything and you get \[\frac{8w^{2}x}{9}\]

OpenStudy (anonymous):

Thank you so much,I need so much help.